OpenStudy (samsan9):

The vertices for the hyperbola (x-2)^2/36-(y+1)^2/64=1 are (2,5) and (2,-7) true or false

OpenStudy (samsan9):

\[\frac{ (x-2)^2 }{ 36 }-\frac{ (y+1)^2 }{ 64 }=1\]